# Classification of Breast Cancer Histopathology Images using a Modified Supervised Contrastive Learning Method

### v1

@@ Abstract
Deep neural networks have reached remarkable achievements in medical image processing tasks, specifically classifying and detecting various diseases. However, when confronted with limited data, these networks face a critical vulnerability, often succumbing to overfitting by excessively memorizing the limited information available. This work addresses the challenge mentioned above by improving the supervised contrastive learning method to reduce the impact of false positives. Unlike most existing methods that rely predominantly on fully supervised learning, our approach leverages the advantages of self-supervised learning in conjunction with employing the available labeled data. We evaluate our method on the BreakHis dataset, which consists of breast cancer histopathology images, and demonstrate an increase in classification accuracy by 1.45% at the image level and 1.42% at the patient level compared to the state-of-the-art method. This improvement corresponds to 93.63% absolute accuracy, highlighting our approach’s effectiveness in leveraging data properties to learn more appropriate representation space. The code implementation of this study is accessible on GitHub [^1] https://github.com/matinamehdizadeh/Breast-Cancer-Detection.
@@ 1 Introduction
Cancer, a leading cause of mortality worldwide, poses a significant threat to public health. Among the various types of cancer, breast cancer is the most prevalent form affecting women, resulting from abnormal cell division in the breast tissue. However, it is essential to note that not all tumors are life-threatening, as they can be either benign or malignant, and detecting a tumor’s malignancy or benign nature is crucial in determining appropriate treatment strategies and ensuring patient well-being.
@@
Medical imaging stands out as one of the most commonly used artificial intelligence applications in healthcare. The research focused on applying machine learning to medical images, which demonstrates significant potential in revolutionizing the diagnostic process of various cancerous diseases by enabling the early detection and accurate characterization of tumors. However, several challenges within the medical field are delaying the widespread utilization of machine-learning capabilities in real-world applications. These challenges include the generalization issues of Deep Neural Networks, collecting diverse data sets, acquiring sufficient labeled data for training, and the risks and high costs associated with errors in medical decision-making. The growing importance of these methods in the medical domain highlights the urgent need to implement effective approaches to overcome these challenges more than ever.
@@
The classification of medical images plays a supportive role in clinical treatment. Histopathology also greatly benefits from the application of machine learning techniques. Leveraging the power of machine learning in diagnosing cancerous images has the potential to revolutionize the field, aiding pathologists in making precise diagnoses and facilitating the treatment process. However, due to the critical nature of this issue and the potential risks, the widespread use of these methods is still limited. As a result, it is crucial to focus on reducing existing errors by examining earlier methods and striving to improve upon them. Supervised methods previously used for image classification have limitations, as they require a considerable amount of labeled data, and collecting them usually requires a high cost. Fortunately, in recent years, self-supervised learning has emerged as a promising approach, presenting its effectiveness in various tasks.
@@
The contributions of this work are focused on overcoming the challenges associated with training networks in the context of cancer image diagnosis. One of the main challenges this study addresses is the scarcity of labeled data in the medical domain, which poses difficulties for supervised learning approaches. To tackle this issue, the proposed method draws inspiration from self-supervised learning techniques and extends them by utilizing labeled data and leveraging similarities between images in the representation space. Additionally, to address the problem of generalization in deep neural networks, various data augmentation techniques specific to histopathology datasets and H&E stains are employed to enhance the robustness and generalization capabilities of the trained models.
@@ 2 Literature Review
Self-supervised learning is an approach that does not need labeled data to train the deep models. It can be divided into two main subcategories: contrastive and non-contrastive. Contrastive learning aims to maximize the similarity between positive pairs and minimize the similarity between negative pairs, e.g., SimCLR [3], SimCLRv2 [4], and PIRL [18] are based on this approach. On the other hand, non-contrastive learning methods work in other ways; for example, they focus only on maximizing the similarity between positive pairs since finding negative pairs can be challenging when working with unlabeled data. Some examples of non-contrastive learning e.g. BYOL [10], DirectPred [25], and SimSiam [5]. Furthermore, a more effective approach to contrastive learning is to utilize the labels associated with the data to guide the choice of negative and positive pairs, where positive pairs consist of samples that belong to the same class, and negative pairs are formed by taking samples from different classes [15].
Research results on medical images using self-supervised methods demonstrate better performance than the supervised ones. For instance, Chhipa et al. [6] introduced a method based on the Efficient-net [23] architecture that utilizes the SimCLR approach for the binary classification of the BreakHis dataset. Jin et al. [14] introduced the HistoSSL architecture, a knowledge distillation model based on non-contrastive learning, which does not rely on a strategy for creating negative pairs. Additionally, Li et al. [16] proposed a combined method of SimCLR and multiple instance learning [8], in which the representations of an image are extracted through self-supervised learning and given as input to a multiple instance learning model. Ciga et al. [7] presented a relatively comprehensive article using the self-supervised SimCLR method on different ResNet [11] architectures, i.e. ResNet-18, ResNet-34, ResNet-50, and ResNet-101. The network was trained in a self-supervised manner on 57 different histopathology datasets, and ultimately, the weights obtained were used for classification, segmentation, and regression tasks on the data.
@@ 3 Method
First, the model is trained on five distinct breast cancer datasets, with the SimCLR technique to establish a solid starting point. We then utilize these pre-trained weights as the initial weights for our representation learning phase. The representation learning phase includes a two-stage supervised contrastive learning. Following this, supervised fine-tuning is accomplished wherein three fully connected layers are added to the end of the encoder for classification task and one fully connected layer is employed for an auxiliary task. Ultimately, four identical architectures are employed in the dataset and trained separately for four magnification factors.
@@ 3.1 Augmentations
The most prevalent way of creating positive pairs is through augmentation. In this study, various methods were utilized for this purpose, including Random Crop, Color Jitter, Gaussian Blur, and Geometric, which are depicted in Figure 1. All of the mentioned augmentations are applied with a fixed probability to create the image pairs, and the resulting pairs can be observed in Figure 2.
@@
By solving this sparse non-negative matrix factorization problem mentioned above, $W$ and $H$ can be estimated [27]. $H$ is the converted matrix, from RGB image to $H$ and $E$ arrays. Subsequently, in HED augmentation, the contrast and brightness of $I_{h}$ and $I_{e}$ are altered as:
@@ 3.2.1 Modified Supervised Contrastive Learning
First, we extract the representation using a modified supervised contrastive loss, which takes the following form.
@@
During the first stage, positive pairs are created using augmentations and image labels. In this context, $i$ is the index of the anchor image. Assuming a set $A$ comprising $12N-$ 1 samples from both images and their corresponding augmented pairs, derived from a batch of size $N$, we define $P$ as the subset of images within set $A$ sharing the same label as the anchor and $Q$ as the subset containing images with different labels from the anchor. The representation of the anchor image is denoted as $z_{i}$, while $z_{p}$ represents one of the positive pairs for $z_{i}$ in the $P$ set. Similarly, $z_{q}$ denotes a negative pair for $z_{i}$ in the $Q$ set. Furthermore, $\alpha$ is the weight assigned to this loss function, which is calculated with respect to the number of samples in both classes to overcome the class imbalance issue. The utilization of the dot product instead of similarity stems from the fact that the data is normalized.
The structure of our presented loss function is similar to Eq. 6. By introducing a coefficient $\lambda$ and the $Q$ set, we have incorporated the effect of the Eq. 7. Additionally, by utilizing augmentations for creating a positive pair for images, we have also integrated the impact of Eq. 5 into our approach. These losses and their individual effectiveness are further analyzed in the result section.
Figure 3: Network architecture. The architecture consists of three training stages, starting with the weights of EfficientNet pre-trained on histopathology images. The pre-training phase involves training on the Breakhis dataset using the modified supervised contrastive learning method. Subsequently, the similarity between the image representations is calculated to correct the positive pairs, followed by retraining with the updated set of pairs in the relaxing phase. The final stage is fine-tuning, which involves an auxiliary and classification task using a supervised approach. Here, the representation of the anchor image is denoted as $z_{i}$, while $z_{p}$ and $z_{q}$ are the representation of a positive pair and a negative pair, respectively. $\hat{y}$ and $\hat{W}=(\hat{w_{1}},\hat{w_{2}},\hat{w_{3}},\hat{w_{4}},\hat{w_{5}},\hat{w_{6}})$ are the predicted label, and conversion matrix with respect to the input image.
@@ 3.2.2 Relaxing Supervised Contrastive Learning
In the relaxing stage, the positive pairs are created like in the previous phase. However, the set $P$ is altered in such a way that if an image in the $P$ set has a similarity score less than 0.5 in the obtained representation space, it is removed from the $P$ set and is no longer considered a positive pair. This means that if, after the model is trained, assuming the $i$-th image is a positive pair for the anchor, which shows negative similarity in the resulted representation space, then it is not a good positive pair. More precisely, it does not have enough similarity even after enforcing it to be similar to the anchor, and so it is removed from the $P$ set. The same technique goes for the $Q$ set: if an image in the $Q$ set has a similarity score of more than 0.5 in the obtained representation space, it is removed from the $Q$ set and is no longer considered a negative pair.
In this stage, we continue to use the same loss function as Eq. 4. However, in contrast to the first stage, $z_{p}$ now represents a positive pair in the $P$ set and is greater than $0.5$ in similarity with the anchor. Similarly, $z_{q}$ denotes a negative pair for $z_{i}$ in the $Q$ set and less than $0.5$ similarity with the anchor. The proposed contrastive learning method is summarized briefly in Algorithm 1.
@@ 3.3.1 Classifier
This architecture, depicted in Figure 3, is designed to perform binary image classification. To accomplish this task, a classification head consisting of three fully connected layers is added at the end of the encoder. This loss function is computed using the following equation.
@@ 3.3.2 Auxiliary Task
The output of this task is a six-element matrix, which predicts the W matrix in Eq. 1. The loss function associated with this branch calculates the difference between the predicted and original matrices. This loss function is computed using the following equation.
@@
The resulting loss value is then fed back to the network with a negative sign to reduce the model’s sensitivity to variations in stain colors. The experiments showed that incorporating the auxiliary task alone does not significantly improve the model’s accuracy. However, to leverage the benefits of this technique, images must first be augmented using HED and then provided as inputs to the encoder. Moreover, during the fine-tuning stage, other data augmentation methods are utilized to further enhance the model’s generalization capabilities and increase the learning process’s robustness.
@@ 4.1 Datasets
    BreakHis [22]. This dataset consists of 7,909 stained microscopic images of breast tumor tissue collected from 82 patients. The images are captured at four different magnification factors, namely, $40X$, $100X$, $200X$, and $400X$. The dataset consists of two primary breast tumor categories, classified as benign or malignant. Specifically, there are 2,480 samples from benign patients and 5,429 from malignant patients. The BreakHis dataset is imbalanced, with significantly more samples in the malignant class. To mitigate the impact of this issue, a weighted loss function was used during all three training stages.
@@
    BACH [1]. This dataset comprises 400 stained microscopic images and 30 whole slide images of breast tumor tissue. The microscopic images have been categorized into normal and benign techniques that have been employed in previous studies, such as those cited in [28]in situ carcinoma and invasive carcinoma, with each category containing 100 images. The purpose of employing the BACH dataset was to evaluate the generalization of our backbone, which was trained using the BreakHis dataset. As the research objective revolves around the binary classification of breast tumors, a few adjustments were made to the labels in the BACH dataset. The normal and benign labels were treated as benign, while the in situ carcinoma and invasive carcinoma labels were considered malignant. Previous studies have employed similar techniques, such as those cited in [28] and [19].
To conduct the experiments, the datasets were split into training, test, and validation sets in a $60\%,20\%,20\%$ ratio, respectively, and the experiments were repeated for 5 folds to ensure accurate results.
@@ 4.2 Experimentation Details
The EfficientNet-B2 [23] is an encoding architecture that is utilized in this study. To ensure a solid starting point, we begin by randomly assigning the Efficient-Net2 weights. Subsequently, we train the model using SimCLR on five distinct breast cancer datasets: andrewjanowczyk epi [13], andrewjanowczyk mitosis [13], andrewjanowczyk nuclei [13], Mitos & Atypia [20], and IDCGrad [2]. In order to evaluate the performance of the proposed model for binary classification of the BreakHis dataset, five previous networks were chosen as benchmarks. Then, their performance was assessed by comparing the Image-Level Accuracy (the average accuracy across all images) and the Patient-Level Accuracy (the average accuracy for individual patients) against the performance of the proposed model.
The network is trained for $200$ epochs in the first two stages. During this process, the temperature of $0.01$, the learning rate $10^{-5}$, and a set of augmentations is utilized. The Adam optimizer is also employed as the optimization function, the batch size is set to $12$, and the $\lambda$ parameter is equal to $2$.
In the next stage of training, i.e., the supervised fine-tuning phase, the four distinct models, each for one magnification factor, are trained for only $20$ epochs. The learning rate is set to $2\times 10^{-5}$, and augmentations are applied to enhance the network’s generalization capabilities. Furthermore, the number of fully connected layers and neurons in each layer are determined through hyperparameter tuning. Additionally, dropout with a probability of $0.5$ is implemented between the fully connected layers, the batch size is 8, and the $\eta$ parameter for the auxiliary task is set to $0.5$.

### v2

@@ Abstract
Deep neural networks have reached remarkable achievements in medical image processing tasks, specifically in classifying and detecting various diseases. However, when confronted with limited data, these networks face a critical vulnerability, often succumbing to overfitting by excessively memorizing the limited information available. This work addresses the challenge mentioned above by improving the supervised contrastive learning method leveraging both image-level labels and domain-specific augmentations to enhance model robustness. This approach integrates self-supervised pre-training with a two-stage supervised contrastive learning strategy. In the first stage, we employ a modified supervised contrastive loss that not only focuses on reducing false negatives but also introduces an elimination effect to address false positives. In the second stage, a relaxing mechanism is introduced that refines positive and negative pairs based on similarity, ensuring that only relevant image representations are aligned. We evaluate our method on the BreakHis dataset, which consists of breast cancer histopathology images, and demonstrate an increase in classification accuracy by 1.45% in the image level, compared to the state-of-the-art method. This improvement corresponds to 93.63% absolute accuracy, highlighting the effectiveness of our approach in leveraging properties of data to learn more appropriate representation space. The code implementation of this study is accessible on GitHub [^1] https://github.com/matinamehdizadeh/Breast-Cancer-Detection.
@@ 1 Introduction
Cancer, a leading cause of mortality worldwide, poses a significant threat to public health. Among the various types of cancer, breast cancer is the most prevalent form affecting women, resulting from abnormal cell division in the breast tissue. However, it is essential to note that not all tumors are life-threatening, as they can be either benign or malignant, and detecting a tumour’s malignancy or benign nature is crucial in determining appropriate treatment strategies and ensuring patient well-being.
@@
Medical imaging stands out as one of the most commonly used applications of artificial intelligence in healthcare. Resent research focused on applying machine learning to medical images demonstrates significant potential in revolutionizing the diagnostic process of various cancerous diseases due to enabling the early detection and accurate characterization of tumors [20, 2]. However, several challenges within the medical field are delaying the widespread utilization of machine-learning capabilities in real-world applications. These challenges include the generalization issues of Deep Neural Networks, collecting diverse data sets, acquiring sufficient labeled data for training, and the risks and high costs associated with errors in medical decision-making. The growing importance of these methods in the medical domain highlights the urgent need to implement effective approaches to overcome these challenges more than ever.
@@
The classification of medical images plays a supportive role in clinical treatment. Histopathology also greatly benefits from the application of machine learning techniques. Leveraging the power of machine learning in diagnosing cancerous images has the potential to revolutionize the field, aiding pathologists to make precise diagnoses and facilitating the treatment process. However, due to the critical nature of this issue and the potential risks, the widespread use of these methods is still limited. As a result, it is crucial to focus on reducing existing errors by examining earlier methods and striving to improve upon them. Supervised methods previously used for image classification have their limitations, as they require a considerable amount of labeled data and collecting them usually needs high cost. Fortunately, in recent years, self-supervised learning has emerged as a promising approach, presenting its effectiveness in various tasks.
@@
The contributions of this work are focused on overcoming the challenges associated with training networks in the context of cancer image diagnosis. One of the main challenges addressed in this study is the scarcity of labeled data in the medical domain, which poses difficulties for supervised learning approaches. To tackle this issue, the proposed method draws inspiration from self-supervised learning techniques and extends them by utilizing labeled data and leveraging similarities between images in the representation space. Additionally, to address the problem of generalization in deep neural networks, various data augmentation techniques specific to histopathology datasets and H&E stains are employed. In the process of diagnosing a carcinogenic tissue sample, pathologists analyze specific features of histopathological images. These features include the morphological characteristics of cellular structures, such as the size and shape of nuclei, the density of cells, and the arrangement of tissues. The color of the sample, influenced by H&E stain, also plays a crucial role in highlighting these features. Hematoxylin stains nuclei blue, while Eosin stains extracellular structures pink, creating a contrast that is essential for differentiating between benign and malignant tissues. However, the consistency of these stains can vary due to differences in reagent concentrations, exposure to light, and tissue scanning techniques, which may lead to color variations across different laboratories. Hence, the augmentation techniques help to enhance the robustness and generalization capabilities of the trained models [22].
@@ 2 Literature Review
Self-supervised learning is an approach that does not need labeled data to train the deep models. It can be divided into two main subcategories: contrastive and non-contrastive. Contrastive learning aims to maximize the similarity between positive pairs and minimize the similarity between negative pairs, e.g., SimCLR [4], SimCLRv2 [5], and PIRL [23] are based on this approach. On the other hand, non-contrastive learning methods works in other ways, for example, they focus only on maximizing the similarity between positive pairs since finding negative pairs can be challenging when working with unlabeled data. Some examples of non-contrastive learning e.g. BYOL [12], DirectPred [34], and SimSiam [6]. Furthermore, a more effective approach to contrastive learning is to utilize the labels associated with the data to guide the choice of negative and positive pairs, where positive pairs consist of samples that belong to the same class and negative pairs are formed by taking samples from different classes [18].
Results of research conducted on medical images using self-supervised methods demonstrate better performance compared to the supervised ones. For instance, Spathis et al. [29], using electrocardiogram data, illustrated that clever transformations of input data with self-supervised approaches can perform well even without extensive annotations. Peng et al. [25] used Information Invariant Clustering to develop local representations without relying on traditional contrastive learning, advancing self-supervised learning for medical image segmentation tasks. Their method introduces a regularized mutual information objective and a boundary-aware loss, enhancing model interpretability. Chhipa et al. [7] introduced a method based on the Efficient-net [32] architecture that utilizes the SimCLR approach for the binary classification of the BreakHis dataset. Jin et al. [16] introduced the HistoSSL architecture, a knowledge distillation model based on non-contrastive learning, which does not rely on a strategy for creating negative pairs. Additionally, Li et al. [19] proposed a combined method of SimCLR and multiple instance learning [9], in which the representations of an image are extracted through self-supervised learning and given as input to a multiple instance learning model. Ciga et al. [8] presented a relatively comprehensive article using the self-supervised SimCLR method on different ResNet [13] architectures, i.e. ResNet-18, ResNet-34, ResNet-50, and ResNet-101. The network was trained in a self-supervised manner on 57 different histopathology datasets, and ultimately, the weights obtained were used for classification, segmentation, and regression tasks on the data. Additionally, Ghesu et al. [10] demonstrated the utility of contrastive learning and online feature clustering on 100 million medical images from diverse modalities, validating their approach on multiple medical problems. Wang et al. [36] proposed the Pyramid-based Local Wavelet Transformer model, which enhances self-supervised learning by integrating a pyramid-based transformer into the encoder of Masked Autoencoders (MAE). This approach is designed to effectively capture both local and global features in histopathological images, improving the representation learning process in self-supervised settings.
@@ 3 Method
First, the model is trained on five distinct breast cancer datasets, with the SimCLR technique to establish a solid starting point. We then utilize these pre-trained weights as the initial weights for our representation learning phase. The representation learning phase includes a two-stage supervised contrastive learning. Following this, supervised fine-tuning is accomplished wherein three fully connected layers are added to the end of the encoder for classification task and one fully connected layer is employed for an auxiliary task. Ultimately, four identical architectures are employed for four magnification factors in the dataset and trained separately.
@@ 3.1 Augmentations
The most prevalent way of creating positive pairs is through augmentation. In this study, various methods were utilized for this purpose, including Random Crop, Color Jitter, Gaussian Blur, and Geometric which are depicted in Figure 1. All of the mentioned augmentations are applied with a fixed probability to create the image pairs, and the resulting pairs can be observed in Figure 2.
@@
By solving this sparse non-negative matrix factorization problem mentioned above, $W$ and $H$ can be estimated [38]. $H$ is the converted matrix, from RGB image to $H$ and $E$ arrays. Subsequently, in HED augmentation, the contrast and brightness of $I_{h}$, and $I_{e}$ are altered as:
@@ 3.2.1 Modified Supervised Contrastive Learning
First, we extract the representation using a modified supervised contrastive loss which takes the following form.
@@
During the first stage, positive pairs are created using augmentations and image labels. In this context, $i$ is the index of the anchor image. Assuming a set $A$ comprising $12N-$ 1 samples, from both images and their corresponding augmented pairs, derived from a batch of size $N$, we define $P$ as the subset of images within set $A$ sharing the same label as the anchor, and $Q$ as the subset containing images with different labels from the anchor. The representation of anchor image is denoted as $z_{i}$, while $z_{p}$ represents one of the positive pairs for $z_{i}$ in the $P$ set. Similarly, $z_{q}$ denotes a negative pair for $z_{i}$ in the $Q$ set. Furthermore, $\alpha$ is the weight assigned to this loss function, which is calculated with respect to the number of samples in both classes to overcome the class imbalance issue. The utilization of the dot product instead of similarity stems from the fact that the data is normalized.
Figure 3: Network architecture. The architecture consists of three training stages, starting with the weights of EfficientNet pre-trained on histopathology images. The pre-training phase involves training on the Breakhis dataset using the modified supervised contrastive learning method. Subsequently, the similarity between the image representations is calculated to correct the positive pairs, followed by a retraining with the updated set of pairs in the relaxing phase. The final stage is the fine-tuning witch involves an auxiliary and classification task using a supervised approach. Here, the representation of the anchor image is denoted as $z_{i}$, while $z_{p}$, and $z_{q}$ are the representation of a positive pair and a negative pair respectively. $\hat{y}$ and $\hat{W}=(\hat{w_{1}},\hat{w_{2}},\hat{w_{3}},\hat{w_{4}},\hat{w_{5}},\hat{w_{6}})$ are the predicted label, and conversion matrix with respect to the input image.
@@ 3.2.2 Relaxing Supervised Contrastive Learning
In the relaxing stage, the positive pairs are created the same as in the previous phase. However, the set $P$ is altered in such a way that if an image in the $P$ set has a similarity score less than 0.5 in the obtained representation space, it is removed from the $P$ set and is no longer considered as a positive pair. This means that if, after the model is trained, assuming the $i$-th image is a positive pair for the anchor, which shows negative similarity in the resulted representation space, then it is not a good positive pair. More precisely, it does not have enough similarity even after enforcing it to be similar to the anchor, and so it is removed from the $P$ set. The same technique goes for the $Q$ set, which is if an image in the $Q$ set has a similarity score more than 0.5 in the obtained representation space, it is removed from the $Q$ set and is no longer considered as a negative pair.
In this stage, we continue to use the same loss funcion as Eq. 4. However, in contrast to the first stage, $z_{p}$ now represents a positive pair in the $P$ set, and greater than $0.5$ similarity with the anchor. Similarly, $z_{q}$ denotes as a negative pair for $z_{i}$ in the $Q$ set, and less than $0.5$ similarity with the anchor. The proposed contrastive learning method is summarized briefly in Algorithm 1.
@@ 3.3.1 Classifier
This architecture, depicted in Figure 3, is designed to perform binary image classification. To accomplish this task, a classification head consist of three fully connected layers, are added at the end of the encoder. This loss function is computed using the following equation.
@@ 3.3.2 Auxiliary Task
The output of this task is a six-element matrix, which predicts the W matrix in Eq. 1. The loss function associated with this branch calculates the difference between the predicted and original matrices: This loss function is computed using the following equation.
@@
The resulting loss value is then fed back to the network with a negative sign to reduce the model’s sensitivity to variations in stain colors. The experiments showed that incorporating the auxiliary task alone does not significantly improve the model’s accuracy. However, to leverage the benefits of this technique, images must first be augmented using HED and then provided as inputs to the encoder. Moreover, during the fine-tuning stage, other data augmentation methods are utilized to enhance the model’s generalization capabilities further and increase the robustness of the learning process.
@@ 4.1 Datasets
    BreakHis [28]. This dataset consists of 7,909 stained microscopic images of breast tumor tissue collected from 82 patients. The images are captured at four different magnification factors, namely, $40X$, $100X$, $200X$, and $400X$. The dataset consists of two primary breast tumour categories, classified as benign or malignant. Specifically, there are 2,480 samples from benign patients and 5,429 from malignant patients. The BreakHis dataset is imbalanced, with significantly more samples in the malignant class. To mitigate the impact of this issue, a weighted loss function was used during all three training stages.
@@
    BACH [1]. This dataset comprises 400 stained microscopic images and 30 whole slide images of breast tumor tissue. The microscopic images have been categorized into normal, benign, in situ carcinoma, and invasive carcinoma, with each category containing 100 images. The purpose of employing the BACH dataset was to evaluate the generalization of our backbone, which was trained using the BreakHis dataset. As the research objective revolves around the binary classification of breast tumors, a few adjustments were made to the labels in the BACH dataset. The normal and benign labels were treated as benign, while the in situ carcinoma and invasive carcinoma labels were considered malignant. Similar technique have been employed in previous studies, such as those cited in [39], and [24].
To conduct the experiments, the datasets was split into training, test, and validation sets in a $60\%,20\%,20\%$ ratio, respectively, and the experiments were repeated for 5 folds to ensure accurate results.
@@ 4.2 Experimentation Details
The EfficientNet-B2 [32] is an encoding architecture that is utilized in this study. To ensure a solid starting point, we begin by randomly assigning the Efficient-Net2 weights. Subsequently, we train the model using SimCLR on five distinct breast cancer datasets: andrewjanowczyk epi [15], andrewjanowczyk mitosis [15], andrewjanowczyk nuclei [15], Mitos & Atypia [26], and IDCGrad [3]. In order to evaluate the performance of the proposed model for binary classification of the BreakHis dataset, some previous networks were chosen as benchmarks. Then, their performance were assessed by comparing the Image-Level Accuracy (the average accuracy across all images) and the Patient-Level Accuracy (the average accuracy for individual patients) against the performance of the proposed model.
The network is trained for $200$ epochs in the first two stages. During this process, the temperature of $0.01$, the learning rate $10^{-5}$, and a set of augmentations is utilized. Additionally, the Adam optimizer is employed as the optimization function, the batch size is set to $12$, and the $\lambda$ parameter is equal to $2$.
In the next stage of training, i.e. the supervised fine-tuning phase, the four distinct models, each for one magnification factor, are trained for only $20$ epochs. The learning rate is set to $2\times 10^{-5}$, and augmentations are applied to enhance the network’s generalization capabilities. Furthermore, the number of fully connected layers and neurons in each layer are determined through hyperparameter tuning. Additionally, dropout with a probability of $0.5$ is implemented between the fully connected layers, the batch size is 8, and the $\eta$ parameter for the auxiliary task is set to $0.5$.